OpenStudy (anonymous):

determine the formula for the nth term of the following sequence: 2,3,6,11,18

OpenStudy (solomonzelman):

what you are doing here, really quickly: \(\large\color{slate}{ 2\color{red}{+1}=3 }\) \(\large\color{slate}{3\color{red}{+3}=6 }\) \(\large\color{slate}{6\color{red}{+5}=11 }\) \(\large\color{slate}{11\color{red}{+7}=18 }\) and on...

OpenStudy (anonymous):

so is that the answer? @SolomonZelman

OpenStudy (solomonzelman):

no,

OpenStudy (anonymous):

 it wouldnt be broken down like this ? @SolomonZelman

OpenStudy (solomonzelman):

you need the formula for the nth term though, \(\large\color{slate}{2\color{red}{+1+(0\times 2)}=3 }\) \(\large\color{slate}{3\color{red}{+1+(1\times 2)}=6 }\) \(\large\color{slate}{6\color{red}{+1+(2\times 2)}=11 }\)

OpenStudy (solomonzelman):

see what I am trying to say?